The Discovery of the Top Quark

Pekka Sinervo University of Toronto

Overview

• Quarks and the Standard Model • The Role of the Top Quark

• The Hunt

• The Discovery

• What We Know About Top Today • Prognostications

Structure of Matter

Some “high school” science

• Matter is made up of atoms

• There are about 109 different species!

• Atom itself has structure:

– consists of small hard “core” (nucleus) ∗ nucleus made of protons and neutrons

– large cloud of electrons

• We can describe all atoms using 3 particle types 1. electrons (e^{−})

2. protons (p)

3. neutrons (n)

But there are hundreds of particles like p and n’s

• Hadrons (π, K, Δ, N^{∗}, ϕ, J/ψ, ...)

• Leptons (e^{−}, μ^{−}, τ^{−}, ν’s)

The Standard Model

The ingredients of the SM are

• 12 fermions – 6 quarks (u, d, s, c, b, t)

– 6 leptons (e^{−}, ν_{e}, μ, ν_{μ}, τ, ν_{τ})

• 3 electroweak bosons (W^{+} , W^{−}, Z^{∘}, γ)

• 8 gluons to describe the strong force

• ?? to describe gravity

Enormously successful paradigm!

• Predicts rich phenomenology

• Accurate! (eg. δ_{M_{z}}/M_{z} ∼ 8 × 10^{−5})

Predicted existence of top quark in late 1970’s

• Higgs boson the only other “essential” missing element

The Quark Picture

“Quarks R Us”

• Up-Up-Down ↔ proton

• Up-Down-Down ↔ neutron

Heavier quarks are unstable

• Decay by emitting a W^{+} or W^{−} boson c → W^{+} s b → W^{−}c

• Lifetimes are τ ∼ 10^{−12} s

All sorts of questions:

• Why are there so many flavours of quarks?

• Why do they have their unique properties?

• What forces act on them?

Why a Top Quark?

In 1977, Lederman et al. discovered a new particle • Mass of 9.54 ± 0.04 GeV/c^{2}

• Decayed to μ^{+} μ^{−} pairs ⇒ A fifth quark

– e^{+} e^{−} production rate ⇒ Q = −1/3

Studies implied that the b quark had a missing partner

• “Flavour-changing neutral current” decays didn’t occur – decays like b → μ^{+} μ^{−}d not observed

• e^{+} e^{−} → bb̄ scattering suggested an SU(2) state ⇒ A sixth quark, known as “top” or “truth”

– Would have Q = +2/3

The Top Quark in the Standard Model

The SM “needs” a heavy top quark

• Loop effects in SM driven by M_{top} – W/Z boson mass ratio

– b − b̄ mixing

However, exact value of M_{top} not constrained

• In 1983, knew it had to be large (> ∼ 20 GeV/c^{2})

• Searches limited by collider centre-of-mass energies

SM Top Quark Phenomenology

Top quarks decay via the weak force

• Predict ∼ 100% t → W^{+} b

• For heavy top (M_{top} > ∼ 100 − 120 GeV/c^{2}), decays before it hadronises

Can be viewed as “pure” β decay

Γ_{top} = \frac{G_{F} M_{t}^{3}}{8√2π} (1 − \frac{M_{W}^{2}}{M_{t}^{2}} )^{2} (1 + 2\frac{M_{W}^{2}}{M_{t}^{2}} )
∼ 175 MeV (\frac{M_{t}}{M_{W}} )^{3} ∼ 2 GeV

• Toponium (tt̄) doesn’t have time to form

• T mesons (tq̄) ∼ degenerate (if they had time...)

Decay dominated by “longitudinal W^{+} ”

• Means the W^{+} prefers helicity 0

• Result of large top and W^{+} mass

The Hunt for the Top Quark

SM didn’t tell us the top quark mass

• Search for top quark production in e^{+} e^{−} and pp̄ – Rate decreases sharply with M_{top}

– Production and decay signature change with mass ∗ pp̄ → W → tb̄

∗ e^{+} e^{−} → tt̄

∗ pp̄ → tt̄ → W^{+} b W^{−}b̄

UA1 experiment at CERN had one “false call”

• Observed a handful of W → tb̄ candidates in 1983 – Turned into a limit of M_{top} > 42 GeV/c^{2} by 1985

• A search by CDF in 1988-89 found M_{top} > 77 GeV/c^{2}

By 1992, had new measurements of M_{Z} and M_{W}

• SM predicted top quark with M_{top} > ∼ 140 GeV/c^{2}

• Only place to look was in high-energy pp̄ collisions • Two experiments

– D0 / and Collider Detector at Fermilab (CDF)

The CDF Detector

Designed to record high-energy p p collisions

• A 5000-ton cylindrical device

• Can create an “image” of the collision

• Originally built in 1982-86

• Seen it’s third (and final?) upgrade

Reconstructing Top

Since top is pair-produced and β decays

• Expect to see 6 final state particles

Life is normally not that easy – in reality:

Identifying Top Decays in pp̄ Collisions

In order to identify top candidates

• Only 1 collision in ∼ 10^{10} expected to produce a top quark • Search for W → lν_{l} candidate

– Electron or muon with P_{T} > 20 GeV/c

– Missing transverse energy E̸_{T} > 20 GeV

• Search for at least one b jet – Most effective tool is to find a secondary vertex

• Some evidence of other W in event – Either two more jets or another lepton

Tagging B Quarks

Must be able to “see” b quarks when searching for top

• Secondary vertex tagging primary technique (SECVTX) – Identify tracks with large impact parameter

– Search for 2 or 3-track combinations with large L_{xy}

• Can tag ∼ 40% of top quark events

Strategy for Detection

Identify events with a W^{+} → l^{+} ν_{l} signature

• Require evidence of a second W boson – W → e^{+} ν_{e} or μ^{+} ν_{μ} – W → q′q̄

• At least one more jet “tagged” as a b quark jet

Look for an excess of such events

Discovery of Top

Hints of top production appeared as early as 1993

• CDF published a 61 page PRD in Sep 1994

• Gave “cagey” label to the reports – “Evidence for top quark production...”

• D0 / ’s analysis not as definitive

With a factor of two more data, result was unequivocal

• April 1995, both collaborations announced discovery

• Confirmed SM predictions – M_{top} ∼ 170 − 190 GeV/c^{2} • Gave us a “free quark” system to study

– Decays before combining with other quarks – No complications associated with QCD force

Convincing “Smoking Gun” Events

2001 CAP UG Lectures

Status of Collider Experiments

CDF and D0 / collected ∼ 110 pb^{−1} of data in 92-96

• 1.8 TeV pp̄ collisions

• Collected events with – High P_{T} leptons (P_{T} > ∼ 10 GeV/c)

– E̸_{T} > 25 GeV

– High P_{T} jets (P_{T} > ∼ 100 GeV/c)

Studies have strengthened top quark interpretation

• Different techniques used to look at the same data – Kinematic distributions

– Neural network approaches

• Verified that all decay modes appear to be OK – See reasonable signals in τ lepton final states – All hadronic modes are also understood

The Status of Top Today

CDF has now detected top in virtually all decay modes

• Lepton + Jets mode (for mass analysis) – lepton P_{T} > 20 GeV/c and E̸_{T} > 20 GeV

– ≥ 3 jets, E_{T} > 15 GeV and |η| < 2

– Require 1 b-tagged jet OR a 4th jet with E_{T} > 15 ⇒ 76 candidates (expect 31 ± 7 bkg)

• Dileptons – 2 leptons with P_{T} > 20 GeV/c

– Remove Drell-Yan and Z^{∘} background ⇒ 9 candidates (2.4 ± 0.5 events bkg)

• All hadronic decays – between 5 and 8 jets (E_{T} > 15 GeV and |η| < 2.0)

– ∑ E_{T} > 300 GeV + additional kinematic cuts ⇒ Observe 187 b-tagged jets (142 ± 11 expected bkg)

What are the Interesting Top Properties?

• Top Quark Mass – SM constrains it to M_{W} /M_{Z}

– Important consistency check of theory

• Top Quark Production Cross Section – Predicted by QCD with less uncertainties than σ_{b} – Look for single top production

• Top Quark Decay Rates – Measured t → W^{+} b/t → W^{+} q

– Look for t → H^{+} b

• Top Quark Decay Kinematics – SM expects decay to be polarized

– Have made first measurement of polarization

• Single Top Production

Top Quark Mass

Measurements now limited by statistical uncertainties

• Overall systematic uncertainty is ∼ 5 GeV/c^{2}

– Dominated by uncertainties in soft gluon contributions

and energy scale

• Run I statistics limit systematic uncertainty

The final Run I values are

M_{top} = 176.1 ± 6.6 GeV c^{2} (CDF)
M_{top} = 172.1 ± 7.1 GeV c^{2} (D0 / )
M_{top} = 174.3 ± 5.1 GeV c^{2} (RunI)

Latest W boson mass results

M_{W} = 80.433 ± 0.079 GeV c^{2} CDF
M_{W} = 80.482 ± 0.091 GeV c^{2} D0
M_{W} = 80.427 ± 0.046 GeV c^{2} LEP2

Top Quark Production Cross Section

The cross section obtained by a fit to

• Lepton + jets, dileptons and all hadronic event rates

• Depends on M_{top} because of acceptance

• Recent improvement is more accurate b-tagging efficiency

σ_{t t} = 6.5_{−1.4}^{+1.7} pb

• Theory predictions range from 4.7 to 5.5 pb

Comparison with D0 /

Measurements by D0 / also quite consistent

Run II Forecasts

Run II has started this spring at √s = 2.0 TeV

• CDF had an “engineering run” (partial SVX II) last fall • Be taking physics data by July 2001

Will be worth the wait

• Begin with ∼ 1 fb^{−1}/year per detector

• Finish up with ∼ 5 fb^{−1}/year

• Literal “Top Factory” with – 300 dilepton events

– 3000 tagged lepton+jet events – 1500 events in “mass sample”

Top Physics Goals in Run II

Mass analysis uncertainties scale with statistics

• Can expect δM_{t} ∼ 3.0˜ GeV/c^{2}

• With δM_{W} ∼ 40 MeV/c^{2} – Begins to hint at the Higgs boson mass – Less than ∼ 210 GeV/c^{2}?

\frac{δM_{Higgs}}{M_{Higgs}} ∼ 40%

• Powerful “pointer” for Higgs searches

Measure angular correlations and decay kinematics

• Tell us about the electroweak force

• Polarization of W^{+} will be measurable

• Spin correlations (expect 90%) also detectable

Single top production

• Measures the “width” of the top quark

• Measure σ_{t} with 10% uncertainty

• Measure V_{tb} with 12% uncertainty

Expected Run IIa Result

2001 CAP UG Lectures

Summary

The debut of top is coming to close

• Have seen that it looks much like a SM fermion

• But top mass scale is enigmatic

All this sets the stage for next set of experiments

• Run for 2-4 years, with goal of ∼ 2 − 15 fb^{−1} ⇒ Increase top quark sample ×50

By the time next generation collider turns on,

• Top will be pretty well-studied

– Reveal some property that points toward a more complete theory?

– Will at least be an important background

• Will be able to reconstruct millions at the LHC!

– It will become the “high-statistics” probe of the next energy scale